OpenStudy (anonymous):

PLEASSSSE SOLVE: solve: -5y+3x=-16 10y+4x=62

OpenStudy (anonymous):

3x=-16+5y x= -16/3 + 5/3y Then plug x into 10y+4x=62 to get the answer for y. Once you have the answer for y, plug that number 3x=-16+5y and get the answer for x.